# Simultaneously Constraining the Neutron Star Equation of State and Mass Distribution through Multimessenger Observations and Nuclear Benchmarks

### v2

@@ I Introduction
The study of neutron stars (NSs) offer unique insights into the behavior of matter under extreme conditions, where the densities are greater than those found in atomic nuclei. To determine the internal composition and structure of a NS one needs to know, apart from the neutron star mass, the equation of state (EOS) that governs matter at supra-nuclear densities Lattimer and Prakash (2016); Oertel et al. (2017); Baym et al. (2018). The EOS, which encapsulates the relationship between pressure and density within the star, is crucial for predicting neutron star key characteristics such as radii, moments of inertia, and tidal deformabilities Hinderer (2008), as well as for the dynamics and final fate of a neutron star merger Baiotti (2019); Shibata and Hotokezaka (2019); Bernuzzi (2020). However, these characteristics are not solely functions of the EOS; they also depend intricately on the star’s mass. Consequently, the accurate determination of NS properties necessitates a simultaneous inference of both mass and EOS.
@@
This article is organized as follows: In Sec II, we provide a description of the construction of the EOS parameterization. We discuss the hybrid parameterization approach that combines nuclear empirical parameters at low densities with piecewise polytropic (PP) models at higher densities. Sec III delves into the multi-messenger observations and the theoretical models employed to constrain the EOS. This section also outlines the hierarchical Bayesian framework used to integrate various datasets. In Sec IV, we employ Gaussian Mixture Models to robustly analyze multi-dimensional observational data from gravitational wave detections, X-ray observations, and empirical EOS parameters, optimizing model selection to capture the complex dependencies within these datasets. Sec V presents the results obtained from the combined analysis of theoretical predictions, observational data, and experimental constraints. We detail the inferred properties of NSs, including the symmetry energy parameters, radius, and tidal deformability. Section VI addresses the various systematic uncertainties that can affect the inference of the equation of state (EOS), highlighting their impact on the accuracy and reliability of EOS determinations. Finally, Sec VII summarizes our findings and discusses their implications for future research, highlighting the synergy between different types of data in enhancing our understanding of NS interiors.
@@ II.3 Hybrid nuclear+PP parameterization:
Density between $n_{\rm match}$ to $1n_{0}\leq n_{1}\leq 2n_{0}$: Chiral effective field theory constraints are reliable up to approximately twice the nuclear saturation density. However, the precise breakdown scale of $\chi$EFT remains uncertain. Recently, Tiwari et al. Tiwari et al. (2023) conducted a Bayesian analysis that combined astrophysical observations with $\chi$EFT constraints, employing two different transition densities, $1.5n_{0}$ and $2n_{0}$. Their study revealed that adopting $2n_{0}$ as the breakdown density yields tighter posteriors. Given the current arbitrariness in selecting the breakdown density, we marginalize over it within the range from $n_{0}$ to $2n_{0}$. To incorporate this, we introduce a free parameter $n_{1}$ into our model which represents the $\chi$EFT breakdown density. In the density range below $n_{1}$, we use a nuclear empirical parameterization described in Sec. II.1. The uncertainties in the lower-order empirical parameters are small and well-constrained Brown and Schwenk (2014); Tsang et al. (2019). Consequently, we fix the lowest-order parameters, such as the energy per nucleon at saturation density, $e_{0}(n_{0})=-15.9~{}\mathrm{MeV}$, the incompressibility, $K_{0}=240~{}\mathrm{MeV}$, and the symmetry energy, $e_{\rm sym}=31.7~{}\mathrm{MeV}$. This leaves the slope parameter $L$ and the curvature parameter $K_{\rm sym}$ of the symmetry energy as the two remaining free parameters in the model.
@@
In total there are 8 free parameters ($\theta_{\epsilon}$) in hybrid+PP EOS parameterization: 2 from nuclear empirical parameterization ($L$ and $K_{\rm sym}$) and 6 from PP parameterization ($n_{1}$, $n_{2}$, $n_{3}$, $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$). For each EOS parameters, we assume a uniform prior and their ranges are shown in Table 1. We have taken broader ranges for all the parameters to avoid such a situation where priors could rail against the posteriors. Finally, it is essential to ensure that the parameterized EOS adheres to fundamental physical principles, specifically causality and monotonicity of pressure with respect to density. Finally, it is crucial to ensure that the parameterized EOS complies with fundamental physical principles, specifically causality and the monotonicity of pressure with respect to density. The speed of sound within the neutron star (NS) must not exceed the speed of light to adhere to the principles of special relativity. Mathematically, this condition is expressed as:
@@
where $S$ denotes the entropy, and the derivative is computed at constant entropy. The pressure must monotonically increase with density to ensure hydrostatic stability within the NS (“Le Chatelier’s principle”, see e.g. Glendenning (1997)):
@@ II.4 Mass distribution model
Previous studies Ozel et al. (2012); Antoniadis et al. (2016); Alsing et al. (2018); Farrow et al. (2019); Farr and Chatziioannou (2020); Shao et al. (2020) indicate that the galactic population of NS masses follow a double Gaussian distribution. However there is no a priori reason to believe that it would be same in other galaxies too. NSs detected through GWs may follow a different distribution than the galactic mass distribution Landry and Read (2021). Nevertheless, given the small number of NSs detected through GWs, a wrong choice of population will have negligible effect on the EOS inference. Therefore we assume for simplicity that all NSs observed in the universe follow a bimodal mass distribution:
@@
We fix $M_{\rm min}=0.9\,M_{\odot}$ to reduce the dimensionality of the parameter space. We do not consider the possibility of low-mass NSs, which is theoretically disfavoured by the early evolutionary stages of a NS Strobel et al. (1999) and is consistent with the minimum remnant masses predicted by core-collapse supernova simulations  Janka et al. (2008); Fischer et al. (2010); Radice et al. (2017); Suwa et al. (2018). $M_{\mathrm{max}}$ is the maximum mass for that particular EOS and the term $U(M|M_{\rm min},M_{\rm max})$ is expressed as follows:
@@ III.2 NS mass Likelihood
Alsing et al. (2018) use three classes of pulsar mass measurements used. The first consists of well-constrained pulsar mass measurements, where the likelihood function is assumed to be (proportional to) a Gaussian:
@@ III.3 GW Likelihood
The advanced LIGO and Virgo detectors have detected GWs from two BNS events named GW170817 and GW190425. Each event’s analysis resulted in a posterior distribution for binary parameters, including component masses $M_{1}$ and $M_{2}$ and tidal deformabilities $\Lambda_{1}$ and $\Lambda_{2}$. Table 2 summarizes key properties of these events. We quote the chirp mass ${\cal{M}}$, mass ratio $q$, and the tidal parameter $\tilde{\Lambda}$ Favata (2014), which is a mass-weighted combination of $\Lambda_{1}$ and $\Lambda_{2}$. It is the best-measured tidal parameter for these events and the only one expected to be measurable with current detector sensitivities Wade et al. (2014). In our analysis, we utilize publicly available posterior samples from LIGO Scientific Collaboration (2018) and LIGO Scientific Collaboration (2020). The posteriors are reported with respect to a prior uniform in $\Lambda_{1}$, $\Lambda_{2}$, and the component masses in the detector frame, with the condition $M_{2}\leq M_{1}$ Abbott et al. (2019, 2020). Mass and tidal deformability measurements from both NSs in a BNS merger serve as macroscopic inputs in our inference of the EOS. The GW likelihood is expressed as:
@@ III.6 Neutron skin measurement likelihood
PREX-II collaboration has reported Adhikari et al. (2021) the value of neutron skin thickness of ${}^{208}{\rm Pb}$ to be, $R_{\rm skin}^{208}=0.29\pm 0.07$ fm (mean and $1\sigma$ standard deviation). Ref. Reed et al. (2021) have argued, this result implies the value of $L$ to be $106\pm 37$ MeV. They also deduce a larger value of the NS radius as there exists a correlation between $R_{1.4}$ and $L$ based on relativistic mean field calculations. However, Ref. Biswas (2021) argued that the measurement uncertainty in $R_{\rm skin}^{208}$ by PREX-II is rather broad. Moreover, the measurement uncertainty deduced after the addition of astrophysical observations is dominated by the latter. Combined astrophysical observations and PREX-II data yield the value of empirical parameter $L=69^{+21}_{-19}$ MeV, $R_{\rm skin}^{208}=0.20_{-0.05}^{+0.05}$ fm, and radius of a $1.4M_{\odot}$ ($R_{1.4})=12.70_{-0.54}^{+0.42}$ km at $1\sigma$ credible interval (CI). Nevertheless, a better measurement of $R_{\rm skin}^{208}$ might have a small effect on the radius of low mass NSs, but for the high masses, there will be almost no effect. State-of-the-art $\chi$EFT calculations predict Xu et al. (2020) the value of $R_{\rm skin}$ to be $0.17-0.18$ fm based on Bayesian analyses using mocked data. Therefore, if the high $R_{\rm skin}^{208}$ values continue to persist with lesser uncertainty, it would pose a challenge to the current theoretical understanding of nuclear matter near saturation densitiy. More recently, the CREX collaboration Adhikari et al. (2022) has reported the value of neutron skin thickness of $\rm{}^{48}\!Ca$ to be $R_{\rm skin}^{48}=0.121\pm 0.026$ (exp) ±$\pm 0.024$ (model) fm. They find several models, including the microscopic coupled cluster calculations Hu et al. (2021) are consistent with the combined CREX and PREX-II results at $90\%$ CI, but in tension at $68\%$ CI. Following Ref. Biswas (2021) we estimate the skin thickness $R_{\text{skin}}$ using the universal relation from Viñas et alViñas et al. (2014) connecting $R_{\text{skin}}$ and the empirical parameter $L$: $R_{\text{skin}}^{208}[\text{fm}]=0.101+0.00147\times L[\text{MeV}]$. Then we also use another empirical relation from Ref. Tripathy et al. (2020), $R_{\text{skin}}^{48}=0.0416+0.6169R_{\text{skin}}^{208}$, to obtain the neutron skin thickness of ${}^{48}\text{Ca}$. The likelihood for a given sample point is simply obtained by comparing to the experimentally determined neutron skin thickness with the result predicted from the phenomenological relations set up by our model collection,
@@ IV Gaussian Mixture Model for Multi-Dimensional Observational Data
Figure 7: Posterior distribution of NS masses after adding the constraints coming from astrophysical observations, nuclear theory and experiments.
@@ VI.1 Systematics in Mass-Radius Posterior Distributions of PSR J0030+0451
The interpretation of NICER data for PSR J0030+4510 is highly sensitive to the choice of model employed to describe the pulsar’s hotspot emission. In this work, we compare the impact of two widely used models, ST+PST and PDT-U, on the mass-radius posterior distributions of neutron stars. These models yield notably different mass and radius estimates for PSR J0030+0451, which, in turn, influence the inferred NS equation of state.
The ST+PST model provides a relatively softer posterior distribution in the mass-radius space. In contrast, the PDT-U model produces a significantly stiffer posterior, indicating larger radii for a given mass. The resultant 90% CI mass-radius posterior distributions, combining other astrophysical and nuclear constraints, are shown in Figure 8. A key observation is the limited overlap between the posterior distributions obtained from the two models. This small overlap highlights the substantial influence of model-dependent systematics in constraining the NS EOS. The stiffness of the PDT-U posterior suggests a preference for equations of state that are more resistant to compression, leading to larger NS radii. Conversely, the softer ST+PST posterior aligns with EOS predictions favoring more compact neutron stars.
@@
Figure 8: Mass-radius posterior distributions for neutron stars obtained using the ST+PST and PDT-U models, combining astrophysical and nuclear constraints. The shaded regions represent the 90% credible intervals.

### v3

@@ I Introduction
The study of neutron stars (NSs) offer unique insights into the behavior of matter under extreme conditions, where the densities are greater than those found in atomic nuclei. To determine the internal composition and structure of a NS one needs to know, apart from the neutron star mass, the equation of state (EOS) that governs matter at supra-nuclear densities Lattimer and Prakash (2016); Oertel et al. (2017a); Baym et al. (2018). The EOS, which encapsulates the relationship between pressure and density within the star, is crucial for predicting neutron star key characteristics such as radii, moments of inertia, and tidal deformabilities Hinderer (2008), as well as for the dynamics and final fate of a neutron star merger Baiotti (2019); Shibata and Hotokezaka (2019); Bernuzzi (2020). However, these characteristics are not solely functions of the EOS; they also depend intricately on the star’s mass. Consequently, the accurate determination of NS properties necessitates a simultaneous inference of both mass and EOS.
@@
This article is organized as follows: In Sec II, we provide a description of the construction of the EOS parameterization. We discuss the hybrid parameterization approach that combines nuclear empirical parameters at low densities with piecewise polytropic (PP) models at higher densities. Sec III delves into the multi-messenger observations and the theoretical models employed to constrain the EOS. This section also outlines the hierarchical Bayesian framework used to integrate various datasets. In Sec IV, we employ Gaussian Mixture Models to robustly analyze multi-dimensional observational data from gravitational wave detections, X-ray observations, and empirical EOS parameters, optimizing model selection to capture the complex dependencies within these datasets. Sec V presents the results obtained from the combined analysis of theoretical predictions, observational data, and experimental constraints. We detail the inferred properties of NSs, including the symmetry energy parameters, radius, and tidal deformability. Sec VI addresses the various systematic uncertainties that can affect the inference of the equation of state (EOS), highlighting their impact on the accuracy and reliability of EOS determinations. Finally, Sec VII summarizes our findings and discusses their implications for future research, highlighting the synergy between different types of data in enhancing our understanding of NS interiors.
@@ II.3 Hybrid nuclear+PP parameterization:
Density between $n_{\rm match}$ to $1n_{0}\leq n_{1}\leq 2n_{0}$: Chiral effective field theory constraints are reliable up to approximately twice the nuclear saturation density. However, the precise breakdown scale of $\chi$EFT remains uncertain. Recently, Tiwari et al. Tiwari et al. (2023) conducted a Bayesian analysis that combined astrophysical observations with $\chi$EFT constraints, employing two different transition densities, $1.5n_{0}$ and $2n_{0}$. Their study revealed that adopting $2n_{0}$ as the breakdown density yields tighter posteriors. Given the current arbitrariness in selecting the breakdown density, we marginalize over it within the range from $n_{0}$ to $2n_{0}$. To incorporate this, we introduce a free parameter $n_{1}$ into our model which represents the $\chi$EFT breakdown density. In the density range below $n_{1}$, we use a nuclear empirical parameterization described in Sec. II.1. The uncertainties in the lower-order empirical parameters are small and well-constrained Brown and Schwenk (2014); Tsang et al. (2019). Consequently, we fix the lowest-order parameters, such as the saturation density $n_{0}=0.16$ fm-3 and the energy per nucleon at saturation density, $e_{0}(n_{0})=-15.9~{}\mathrm{MeV}$. A survey of 53 experimental results in 2016 Oertel et al. (2017b) found symmetry energy and compressibility values of $e_{\rm sym}(\rho_{0})=31.7\pm 3.2$ MeV and $K_{0}=240\pm 30$ MeV. While there are larger uncertainties on these parameters, previous subsequent Bayesian analyses Biswas et al. (2021); Biswas (2021); Char and Biswas (2024), which combined multiple astrophysical observations (gravitational waves and X-ray data), found that current observational constraints do not provide significant information on these parameters. Given this, we chose to fix $e_{\rm sym}(\rho_{0})$ and $K_{0}$ at their median values from the survey—31.7 MeV and 240 MeV, respectively—rather than unnecessarily increasing the dimensionality of the EOS parameterization. Since the present data does not favor or constrain a particular range of $e_{\rm sym}$ and $K_{0}$, our results remain robust under reasonable variations of these parameters. This leaves the slope parameter $L$ and the curvature parameter $K_{\rm sym}$ of the symmetry energy as the two remaining free parameters in the model.
@@
In total there are 8 free parameters ($\theta_{\epsilon}$) in hybrid+PP EOS parameterization: 2 from nuclear empirical parameterization ($L$ and $K_{\rm sym}$) and 6 from PP parameterization ($n_{1}$, $n_{2}$, $n_{3}$, $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$). For each EOS parameters, we assume a uniform prior and their ranges are shown in Table 1. We have taken broader ranges for all the parameters to ensure flexibility in the equation of state and to avoid scenarios where priors could strongly constrain the posteriors. This approach allows the data to guide the inference without imposing unnecessary restrictions. Similar strategies have been adopted in previous studies, such as Refs. Hebeler et al. (2013); Raaijmakers et al. (2020). Finally, it is crucial to ensure that the parameterized EOS complies with fundamental physical principles, specifically causality and the monotonicity of pressure with respect to density. The speed of sound within the neutron star (NS) must not exceed the speed of light to adhere to the principles of special relativity. Mathematically, this condition is expressed as:
@@
where the derivative is computed at constant entropy $S$. The pressure must monotonically increase with density to ensure hydrostatic stability within the NS (“Le Chatelier’s principle”, see e.g. Glendenning (1997)):
@@ II.4 Mass distribution model
Previous studies Ozel et al. (2012); Antoniadis et al. (2016); Alsing et al. (2018); Farrow et al. (2019); Farr and Chatziioannou (2020); Shao et al. (2020); Das et al. (2024) indicate that the galactic population of NS masses follow a double Gaussian distribution. However there is no a priori reason to believe that it would be same in other galaxies too. NSs detected through GWs may follow a different distribution than the galactic mass distribution Landry and Read (2021). Nevertheless, given the small number of NSs detected through GWs, a wrong choice of population will have negligible effect on the EOS inference. Therefore we assume for simplicity that all NSs observed in the universe follow a bimodal mass distribution:
@@
We fix $M_{\rm min}=0.9\,M_{\odot}$ to reduce the dimensionality of the parameter space. We do not consider the possibility of low-mass NSs, which is theoretically disfavoured by the early evolutionary stages of a NS Strobel et al. (1999) and is consistent with the minimum remnant masses predicted by core-collapse supernova simulations  Janka et al. (2008); Fischer et al. (2010); Radice et al. (2017); Suwa et al. (2018). $M_{\mathrm{max}}$ is the maximum mass for that particular EOS and the term $U(M|M_{\rm min},M_{\rm max})$ reads:
@@ III.2 NS mass Likelihood
Alsing et al. (2018) use three classes of pulsar mass measurements. The first consists of well-constrained pulsar mass measurements, where the likelihood function is assumed to be (proportional to) a Gaussian:
@@ III.3 GW Likelihood
The advanced LIGO and Virgo detectors have detected GWs from two BNS events named GW170817 and GW190425. Each event’s analysis resulted in a posterior distribution for binary parameters, including component masses $M_{1}$ and $M_{2}$ and tidal deformabilities $\Lambda_{1}$ and $\Lambda_{2}$. Table 2 summarizes key properties of these events. We quote the chirp mass ${\cal{M}}$, mass ratio $q$, and the tidal parameter $\tilde{\Lambda}$ Favata (2014), which is a mass-weighted combination of $\Lambda_{1}$ and $\Lambda_{2}$. It is the best-measured tidal parameter for these events and the only one expected to be measurable with current detector sensitivities Wade et al. (2014). In our analysis, we utilize publicly available posterior samples from LIGO Scientific Collaboration and Virgo Collaboration (2018) and LIGO Scientific Collaboration and Virgo Collaboration (2020). The posteriors are reported with respect to a prior uniform in $\Lambda_{1}$, $\Lambda_{2}$, and the component masses in the detector frame, with the condition $M_{2}\leq M_{1}$ Abbott et al. (2019, 2020). Mass and tidal deformability measurements from both NSs in a BNS merger serve as macroscopic inputs in our inference of the EOS. The GW likelihood is expressed as:
@@ III.6 Neutron skin measurement likelihood
PREX-II collaboration has reported Adhikari et al. (2021) the value of neutron skin thickness of ${}^{208}{\rm Pb}$ to be, $R_{\rm skin}^{208}=0.29\pm 0.07$ fm (mean and $1\sigma$ standard deviation). Ref. Reed et al. (2021) have argued, this result implies the value of $L$ to be $106\pm 37$ MeV. They also deduce a larger value of the NS radius as there exists a correlation between $R_{1.4}$ and $L$ based on relativistic mean field calculations. However, Ref. Biswas (2021) argued that the measurement uncertainty in $R_{\rm skin}^{208}$ by PREX-II is rather broad. Moreover, the measurement uncertainty deduced after the addition of astrophysical observations is dominated by the latter. Combined astrophysical observations and PREX-II data yield the value of empirical parameter $L=69^{+21}_{-19}$ MeV, $R_{\rm skin}^{208}=0.20_{-0.05}^{+0.05}$ fm, and radius of a $1.4M_{\odot}$ ($R_{1.4})=12.70_{-0.54}^{+0.42}$ km at $1\sigma$ credible interval (CI). Nevertheless, a better measurement of $R_{\rm skin}^{208}$ might have a small effect on the radius of low mass NSs, but for the high masses, there will be almost no effect. State-of-the-art $\chi$EFT calculations predict Xu et al. (2020) the value of $R_{\rm skin}$ to be $0.17-0.18$ fm based on Bayesian analyses using mocked data. Therefore, if the high $R_{\rm skin}^{208}$ values continue to persist with lesser uncertainty, it would pose a challenge to the current theoretical understanding of nuclear matter near saturation density. More recently, the CREX collaboration Adhikari et al. (2022) has reported the value of neutron skin thickness of $\rm{}^{48}\!Ca$ to be $R_{\rm skin}^{48}=0.121\pm 0.026$ (exp) ±$\pm 0.024$ (model) fm. They find several models, including the microscopic coupled cluster calculations Hu et al. (2021) are consistent with the combined CREX and PREX-II results at $90\%$ CI, but in tension at $68\%$ CI. Following Ref. Biswas (2021) we estimate the skin thickness $R_{\text{skin}}$ using the universal relation from Viñas et alViñas et al. (2014) connecting $R_{\text{skin}}$ and the empirical parameter $L$: $R_{\text{skin}}^{208}[\text{fm}]=0.101+0.00147\times L[\text{MeV}]$. Then we also use another empirical relation from Ref. Tripathy et al. (2020), $R_{\text{skin}}^{48}=0.0416+0.6169R_{\text{skin}}^{208}$, to obtain the neutron skin thickness of ${}^{48}\text{Ca}$. The likelihood for a given sample point is simply obtained by comparing to the experimentally determined neutron skin thickness with the result predicted from the phenomenological relations set up by our model collection,
@@ IV Gaussian Mixture Model for Multi-Dimensional Observational Data
Figure 6: Evolution of the inferred distribution of crust-core junction density as we add the constraints successively.
@@
Figure 8: Posterior distribution of total NS masses after adding the constraints coming from astrophysical observations, nuclear theory and experiments.
@@ V Results
  • Crust-core junction density: In our study, the crust EOS SLy is modeled using a piecewise polytropic representation, with the corresponding parameters taken from Read et al. Read et al. (2009). The last polytropic index of SLy is applied at a density of $2.62789\times 10^{12}$ g/cm3. The matching density, $n_{\rm match}$, lies between this point and $n_{1}$, and it is determined by ensuring continuity in pressure between the SLy crust and our empirical EOS model. To further examine the impact of $n_{\rm match}$, we have quantified its posterior distribution for each dataset, applying successive constraints. The results are presented in Fig. 6, where we show the evolution of the posterior as additional constraints are incorporated, including gravitational waves (GWs), NICER observations, pulsar mass measurements (PSRs), chiral effective field theory (EFT), perturbative QCD (pQCD), and nuclear experiments such as CREX and PREX-II. From the figure, it is evident that the inclusion of chiral EFT constraints has a significant impact on the inferred transition density, shifting the posterior towards lower values of $n_{\rm match}$. This suggests that the crust-core transition is strongly influenced by nuclear physics constraints at sub-saturation densities.
@@ VI.1 Systematics in Mass-Radius Posterior Distributions of PSR J0030+0451
The reanalysis of NICER data for PSR J0030+4510 conducted by Ref. Vinciguerra et al. (2023) demonstrates a strong dependence on the choice of hotspot emission model. As discussed in Sec. III.4, both the ST+PDT and PDT-U models are capable of reproducing NICER and XMM-Newton data. These models yield notably different mass and radius estimates for PSR J0030+0451, which, in turn, influence the inferred NS equation of state.
The ST+PDT model provides a relatively softer posterior distribution in the mass-radius space. In contrast, the PDT-U model produces a significantly stiffer posterior, indicating larger radii for a given mass. The resultant 90% CI mass-radius posterior distributions, combining other astrophysical and nuclear constraints, are shown in Figure 9. A key observation is the limited overlap between the posterior distributions obtained from the two models. This small overlap highlights the substantial influence of model-dependent systematics in constraining the NS EOS. The stiffness of the PDT-U posterior suggests a preference for equations of state that are more resistant to compression, leading to larger NS radii. Conversely, the softer ST+PDT posterior aligns with EOS predictions favoring more compact neutron stars.
@@
Figure 9: Mass-radius posterior distributions for neutron stars obtained using the ST+PDT and PDT-U models, combining astrophysical and nuclear constraints. The shaded regions represent the 90% credible intervals.